### v1

# Creation of Wavepackets for Quantum Chromodynamics on Quantum Computers

Abstract
One of the most ambitious goals of quantum simulation of quantum field theory is the description of scattering in real time, which would allow not only for computation of scattering amplitudes, but also for studying the collision process step by step. The initial state of such simulation is made of typically two wavepackets moving on top of the vacuum, whose preparation is hard. Some proposals are based on a modified adiabatic transformation of free wavepackets into interacting wavepackets, but this is not suitable for quantum chromodynamics (QCD). Other approaches used in one-dimensional lattice gauge theories rely on an adiabatic transformation from the strong coupling limit, or on ansatzes to create the wavepackets from the full, interacting vacuum. Here we extend a previous work to create wavepackets of a general kind of particle from the vacuum in lattice gauge theories in one and three dimensions, including QCD. The conceptual foundation of this approach is the Haag-Ruelle scattering theory, and the only theoretical limitation is given by massless particles. In the context of digital quantum computation, the wavepacket creation from the vacuum is implemented with the technique known as LCU (linear combination of unitaries). The preparation is performed successfully with a certain success probability, which may represent the main obstacle of our approach.
I Introduction
Quantum computers are expected to have a large impact on the problems that present an exponential improvement with respect to classical computation. Of the few known problems, simulation of complex quantum systems is probably the most prominent. Among these, strongly-correlated many-body systems are very hard to deal with, because traditional approaches fail except in a handful of cases, or for limited classes of questions. Quantum and quantum-inspired approaches are promising solutions to this dilemma [1, 2, 3, 4, 5, 6, 7, 8].
A quantum field theory is essentially a strongly-correlated quantum system with particular requirements on the symmetries it enjoys, namely Poincaré invariance. At a deep level, the correct implementation of this symmetry group also seems to demand gauge invariance to obtain nontrivial dynamics, at least in a three-dimensional space. The most challenging quantum field theory is quantum chromodynamics (QCD), which describes the strong interaction between quarks and gluons. Differently from all the other sectors of the Standard Model of particle physics, a large part of its phenomenology is intrinsically nonperturbative, whence analytical techniques can be used only to set up the right framework for numerical calculations. The gauge group is $SU(3)$, which increases computational costs by several orders of magnitude with respect to a $U(1)$ theory, to make a comparison. Traditional lattice calculations can only address problems amenable to the imaginary-time formalism, such as prediction on particle masses and matrix elements. Nonperturbative studies of processes that depend on real-time evolution is one of the major open problems of particle physics.
For the reasons just mentioned, considerable effort within the quantum-computing community and the high-energy-physics one points towards the ultimate goal of simulating real-time evolution of processes within QCD [9, 10, 11, 12, 13]. The resources required to perform a useful quantum calculation in QCD are several orders of magnitude beyond the present capabilities [14], but promising progress has been made in the last couple of years towards the realization of a large-scale, error-corrected quantum computer, both on the experimental and the theoretical side. Although a full quantum computer is still out of reach, it nonetheless makes sense to analyze all the possibilities to obtain realistic resource estimates, and determine which questions we will be able to address.
Specifically, we consider scattering of particles, a class of experiments which is extremely important for our understanding of fundamental physics. At very high energy, QCD is perturbative, and perturbation theory is well suited to study cross sections and test the theory against experimental results. At low energies, perturbation theory is no longer useful, so it is very challenging to study, for example, pion-pion or proton-proton collisions above the inelastic-scattering threshold. Quantum simulation offers a possibility to study the real-time evolution of these processes, but compared to other applications, the initial state preparation is particularly difficult. In the framework of the Hamiltonian formulation on a spatial lattice, our goal is to prepare two wavepackets of some kind of particles on top of the vacuum. For other approaches that avoid the necessity of wavepackets, see [15, 16]. In most existing approaches, the key step consists of a modified adiabatic transformation to turn wavepackets of the free theory into wavepackets of the interacting theory [17, 18, 19, 20, 21]. Since in the free theory there are no states corresponding to composite particles, this method can only be used for particles associated with elementary degrees of freedom. As a consequence of confinement in QCD, no quarks or gluons appear in the spectrum, and all the particles (hadrons) are composite. Therefore, the works in [17, 18, 19, 20, 21] do not apply to QCD.
There are a few ideas in the literature to circumvent this problem: variational approaches [22], variational approaches combined with adiabatic transformation from the strong-coupling limit [23], and optimized ansatzes [24]. An approach which does not require optimization or ansatzes is [25], based on the Haag-Ruelle scattering theory. In all these methods, wavepackets are prepared directly from the vacuum of the full, interacting theory, which avoids the necessity for adiabatic transformations. In the context of theories with a mass gap, the approach in [25] is fully general, and easily adaptable to different formulations and different particles. Furthermore, its asymptotic efficiency can be analyzed theoretically. On the other hand, the works in [22, 23, 24] have been tested numerically only in simple models in one dimension so far, and it is difficult to perform the analysis of the asymptotical complexity. Generalizations to higher dimensions are possible but not straightforward.
In this work, we show how to apply the method in [25] to lattice gauge theories. As a warm up, we consider $U(1)$, $SU(2)$ and $SU(3)$ theories in one dimension. Subsequently, we address QCD in three dimensions with two flavors of quarks. Not only the one-dimensional theories provide simpler versions of QCD where to make some tests, but also the tools that are used in the one-dimensional case are also used to build up the techniques in the three-dimensional case. As examples, we consider the pions and the nucleons, while other kinds of particles can be treated in a completely analogous way.
We work in the Kogut-Susskind formulation, since this is the choice of preference in many works on quantum simulation of lattice gauge theories in the Hamiltonian formulation. However, our conclusion is that in three dimensions this is not a good choice, because the lattice fermionic degrees of freedom do not correspond with the convenient degrees of freedom in the continuum (Dirac spinors). In addition to the problems already addressed in [26], the Kogut-Susskind formulation has two more disadvantages when we deal with particle scattering: 1) large overheads in the creation of wavepackets with our approach; 2) mixing of flavor symmetries and cubic symmetries, which makes the concepts of spin and isospin even less precise. While the first problem is specific to our method, the second one is not. Applying our ideas to formulations other than the Kogut-Susskind one, like Wilson fermions, is simpler.
The rest of this paper is organized as follows. In section II, we review the approach presented in [25], and we extend it to particles of general spin. In section III we give the lattice operators that are necessary for the creation of wavepackets in one-dimensional lattice gauge theories $U(1)$, $SU(2)$ and $SU(3)$. In section IV we show how to implement the same operators in terms of qubit operators. Finally, in section V we consider QCD in three dimensions. In the appendix we review the generalized superfast encoding of fermions onto qubits, and we also show how to implement odd operators.
II Creation of Wavepackets from the Vacuum
In this section we briefly review the main features of the method proposed in [25], based on the Haag-Ruelle scattering theory, to create wavepackets of particles in an interacting field theory from the vacuum of the theory. The Haag-Ruelle formulation [27, 28] was originally developed in axiomatic quantum field theory and later on formulated for lattice systems [29]. For a review of the formulation in the continuum, see [30].
We can create wavepackets of a certain particle using an operator of the form
$\hat{a}_{\psi}^{\dagger}=\int\!d^{D}x\,\psi(x)\hat{\mathcal{O}}(x),$ (1)
where $D=d+1$ is the number of spacetime dimensions, $x$ is a spacetime coordinate, $\psi$ is a Schwartz function, and $\hat{\mathcal{O}}$ is an interpolating Heisenberg operator for the particle we want to create. The function $\psi$ and the operator $\hat{\mathcal{O}}$ are chosen considering the spectrum and the symmetries of the theory.
For example, suppose there is only one particle of mass $m$ in the spectrum. Then, the joint energy-momentum spectrum is made of three disjoint sectors:
  • the vacuum, corresponding to the origin, $p_{\mu}=0$;
  • the one-particle mass hyperboloid, given by the points such that $p_{\mu}p^{\mu}=m^{2}$;
  • the multi-particle continuum, defined by $p_{\mu}p^{\mu}\geq 4m^{2}$.
The function $\psi$ is chosen as the Fourier transform of a function $\tilde{\psi}$ with support overlapping only with the one-particle mass-hyperboloid, and not with the multi-particle continuum, as depicted in figure 1. The interpolating operator $\hat{\mathcal{O}}$ should carry the same quantum numbers as the particle. More precisely, if $|\alpha\rangle$ is a one-particle state and $|\Omega\rangle$ the vacuum, then the condition for $\hat{\mathcal{O}}$ to be an interpolating operator is
$\langle\alpha|\hat{\mathcal{O}}|\Omega\rangle\neq 0.$ (2)
As a second example, suppose there are two particles of masses $m_{1}$ and $m_{2}$. Our goal is to create wavepackets of a definite kind of particle, and not a mixture of the two. If $m_{2}$ lies halfway between $m_{1}$ and $2m_{1}$, this can be achieved by choosing $\tilde{\psi}$ with support only on the mass hyperboloid of one kind of particle, namely either around $p_{\mu}p^{\mu}=m_{1}^{2}$ or $p_{\mu}p^{\mu}=m_{2}^{2}$. If $m_{2}$ is close to $m_{1}$, close to $2m_{1}$, or is larger than $2m_{1}$, then the operator $\hat{\mathcal{O}}$ must be chosen to be an interpolating operator only for one kind of particle, and not for the other. For example, we could have
$\langle\alpha_{1}|\hat{\mathcal{O}}|\Omega\rangle\neq 0,\qquad\langle\alpha_{2}|\hat{\mathcal{O}}|\Omega\rangle=0,$ (3)
where $|\alpha_{i}\rangle$ is a state of particle $i$. The way to identify particles, and hence appropriate interpolating operators, is through symmetry considerations: the interpolating operator should carry the same quantum numbers as the particle in consideration.
Figure 1: Structure of the joint spectrum of $P_{\mu}$ in a one-dimensional space. The blue line represents the one-particle mass hyperboloid, the red region is the multiparticle continuum, and the green spot represents the support of the function $\tilde{\psi}$.
When we pass from the continuum spacetime to the lattice, we use the same approach. The main difficulty is that on the lattice we typically have only discrete versions of the symmetries in the continuum, and there may be different kinds of particles that are not present in the continuum limit. When we discretize the theory on a space lattice with spacing $a$ to make a quantum simulation, equation (1) becomes
[…]
where $\hat{\mathcal{O}}(\mathbf{x})$ is the Schrödinger representation of $\hat{\mathcal{O}}(x)$, and $x=(t,\mathbf{x})$. The lattice theory needs to be mapped onto a qubit system, with all due truncation and digitization [31, 32]. Also the time integral in equation (4) needs to be approximated by a discrete and truncated summation in steps of $\delta_{t}$,
$\hat{a}_{\psi}^{\dagger}=\sum_{\mathbf{x},t}a^{d}\delta_{t}\psi(t,\mathbf{x})e^{iHt}\hat{\mathcal{O}}(\mathbf{x})e^{-iHt}.$ (5)
Once we know how to implement $\hat{\mathcal{O}}$ on a quantum computer, the creation operator $\hat{a}_{\psi}^{\dagger}$ can be implemented by means of a linear combination of unitaries (LCU) [33]. If $T$ is the size of $\psi$ in the time direction, this requires simulating the evolution for a time $O(T)$.
The interpolator $\hat{\mathcal{O}}(\mathbf{x})$ can be typically written as a LCU itself,
$\hat{\mathcal{O}}(\mathbf{x})=\sum_{j}c_{j}\mathcal{U}_{j}(\mathbf{x}),\qquad\sum_{j}|c_{j}|=C.$ (6)
Overall, $\hat{a}_{\psi}^{\dagger}$ is successfully implemented with probability
$\rho=\bigg{(}\frac{\big{\|}\hat{a}_{\psi}^{\dagger}|\Omega\rangle\big{\|}}{\alpha}\bigg{)}^{2},\qquad\text{where }\alpha=C\sum_{\mathbf{x},t}a^{d}\delta_{t}|\psi(t,\mathbf{x})|.$ (7)
If the operator $\hat{\mathcal{O}}(x)$ is a scalar field, the analysis in [25] shows that the behavior of $\rho$ is
$\rho\sim\frac{Z}{C^{2}}\frac{\delta_{p}^{d}}{\bar{E}},$ (8)
where $\delta_{p}$ is the extension of the support of $\tilde{\psi}$ in the space directions, $\bar{E}$ is the energy of the particle, and $Z$ is the squared renormalization constant connecting $\hat{\mathcal{O}}(\mathbf{x})$ with its continuum normalization, typically logarithmic in the lattice spacing [34].
Generally, $\hat{\mathcal{O}}(x)$ can be a spinorial field (or another kind of field), as is the case for the proton, and therefore the result needs to be adapted. Assume that under a Lorentz transformation in the continuum we have $\hat{\mathcal{O}}_{\alpha}\to\sum_{\beta}S(\Lambda)_{\alpha\beta}\hat{\mathcal{O}}_{\beta}$, and that the operators $\hat{\mathcal{O}}_{\alpha}$ interpolate for a particle of spin $J$ and mass $m$. There are $12J+$ 1 one-particle states with momentum $\mathbf{k}$, $|\mathbf{k},j\rangle$, with $j=-J,-J+1,\dots,J$. Then we have
$\big{\|}\hat{a}_{\alpha,\psi}^{\dagger}|\Omega\rangle\big{\|}^{2}=\int d^{3}k\big{|}\tilde{\psi}(\mathbf{k})\big{|}^{2}\frac{m}{E(\mathbf{k})}\sum_{j}\Big{|}\sum_{\beta}S(\Lambda_{\mathbf{k}})_{\alpha\beta}\sqrt{Z_{\beta}^{j}}\Big{|}^{2},$ (9)
where $E(\mathbf{k})=\sqrt{|\mathbf{k}|^{2}+m^{2}}$, $\tilde{\psi}(\mathbf{k})=\tilde{\psi}(E(\mathbf{k}),\mathbf{k})$, $\sqrt{Z_{\beta}^{j}}=\langle\mathbf{0},j|\hat{\mathcal{O}}_{\beta}(0)|\Omega\rangle$, and $\Lambda_{\mathbf{k}}$ is the boost that sends $\mathbf{k}$ to $\mathbf{0}$. If $\tilde{\psi}$ is narrowly concentrated around $\bar{\mathbf{k}}$, then we roughly have $S(\Lambda_{\mathbf{k}})\sim S(\Lambda_{\bar{\mathbf{k}}})$ in equation (9). Therefore, in equation (8), we can replace $Z$ with
$Z\rightarrow\sum_{j}\Big{|}\sum_{\beta}S(\Lambda_{\bar{\mathbf{k}}})_{\alpha\beta}\sqrt{Z_{\beta}^{j}}\Big{|}^{2}.$ (10)
We consider now the example of a spinorial field in the Dirac representation interpolating for a spin $1/2$ particle with momentum $\bar{\mathbf{k}}=(k_{x},0,0)$. We have
$S(\Lambda_{\bar{\mathbf{k}}})=\begin{pmatrix}\sqrt{\frac{\gamma+1}{2}}\mathbb{1}&\sqrt{\frac{\gamma-1}{2}}\sigma_{x}\\ \sqrt{\frac{\gamma-1}{2}}\sigma_{x}&\sqrt{\frac{\gamma+1}{2}}\mathbb{1}\end{pmatrix},\qquad\gamma=\frac{E(\mathbf{\bar{\mathbf{k}}})}{m}.$ (11)
Then, for $k_{x}\gg m$, we have
$\sum_{\beta}S(\Lambda_{\bar{\mathbf{k}}})_{\alpha\beta}\sqrt{Z_{\beta}^{j}}=\frac{1}{\sqrt{2}}\sqrt{\frac{k_{x}}{m}}\Big{(}\sqrt{Z_{1}^{j}}-\sqrt{Z_{4}^{j}}\Big{)}+\frac{1}{2\sqrt{2}}\sqrt{\frac{m}{k_{x}}}\Big{(}\sqrt{Z_{1}^{j}}+\sqrt{Z_{4}^{j}}\Big{)}+\dots$ (12)
We see that, for a spinorial interpolator, and if $Z_{1}^{j}\neq Z_{4}^{j}$, the term $k_{x}/m$ from this expansion compensates the term $\bar{E}$ in the numerator of equation (8) for highly energetic spin-$1/2$ particles.
III Interpolating Operators in One-dimensional Lattice Gauge Theories
We start by considering the simpler case of gauge theories in one space dimension, which is enough to show all the building blocks that are necessary in the physical case of quantum chromodynamics. The Kogut-Susskind Hamiltonians for lattice gauge theories $U(1)$, $SU(2)$ and $SU(3)$ in one space dimension can be compactly written as
$H=\sum_{x=0}^{N-1}\Big{[}\frac{i}{2a}\big{(}\xi^{\dagger}(x+1)U(x)\xi(x)-\xi^{\dagger}(x)U^{\dagger}(x)\xi(x+1)\big{)}+(-1)^{x}m_{0}\xi^{\dagger}(x)\xi(x)+\frac{ag_{0}^{2}}{2}E(x)^{2}\Big{]}.$ (13)
When referring to the $U(1)$ theory, $\xi(x)$, $U(x)$ and $E(x)$ are one-component fields. When referring to the $SU(2)$ theory, the same symbols mean
$\xi(x)=\begin{pmatrix}\xi_{1}(x)\\ \xi_{2}(x)\end{pmatrix},\qquad U(x)=\begin{pmatrix}U_{11}(x)&U_{12}(x)\\ U_{21}(x)&U_{22}(x)\end{pmatrix},\qquad E(x)^{2}=\sum_{a=1}^{3}E_{\text{L}}^{a}(x)^{2}=\sum_{a=1}^{3}E_{\text{R}}^{a}(x)^{2}.$ (14)
When referring to the $SU(3)$ theory, the symbols mean
$\xi(x)=\begin{pmatrix}\xi_{1}(x)\\ \xi_{2}(x)\\ \xi_{3}(x)\end{pmatrix},\qquad U(x)=\begin{pmatrix}U_{11}(x)&U_{12}(x)&U_{13}(x)\\ U_{21}(x)&U_{22}(x)&U_{23}(x)\\ U_{31}(x)&U_{32}(x)&U_{33}(x)\\ \end{pmatrix},\qquad E(x)^{2}=\sum_{a=1}^{8}E_{\text{L}}^{a}(x)^{2}=\sum_{a=1}^{8}E_{\text{R}}^{a}(x)^{2}.$ (15)
In these formulations we use staggered fermions. A relevant caveat is that only even lattice translations can be interpreted as space translations in the continuum. Translations that involve an odd number of lattice units in one or more directions do not correspond to space translations in the continuum, but to other transformations that are not necessarily symmetries of the theory. Hence, the summation over $\mathbf{x}$ in equation (4) should be in steps of two, as in
[…]
III.1 $U(1)$ Gauge Theory
For the $U(1)$ gauge theory, also known as the Schwinger model, the fields satisfy the following anticommutation and commutation relations:
$\{\xi(x),\xi^{\dagger}(y)\}=\delta_{xy},\qquad[U(x),E(y)]=\delta_{xy}U(x).$ (17)
The local Hilbert space of the gauge field can be expressed in the electric-field basis {$\{|\mathcal{E}\rangle:\mathcal{E}\in\mathbb{Z}\}$, in which we have
$E|\mathcal{E}\rangle=\mathcal{E}|\mathcal{E}\rangle,\qquad U|\mathcal{E}\rangle=|\mathcal{E}-1\rangle.$ (18)
In this model there are three stable mesons in the continuum limit, with associated interpolating operators
$\bar{\psi}\psi,\qquad\psi^{\dagger}\gamma_{5}\psi,\qquad i\bar{\psi}\gamma_{5}\psi.$ (19)
In the staggered-fermion formulation these operators translate respectively to
$\Phi_{1}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)\xi(x)-\xi^{\dagger}(x+1)\xi(x+1)\Big{]},$ (20)
$\Phi_{2}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)U^{\dagger}(x)\xi(x+1)+\xi^{\dagger}(x+1)U(x)\xi(x)\Big{]},$ (21)
$\Phi_{3}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)U^{\dagger}(x)\xi(x+1)-\xi^{\dagger}(x+1)U(x)\xi(x)\Big{]},$ (22)
where we have dropped the factor $i$ in $\Phi_{3}$ as it is irrelevant for our purposes.
III.2 $SU(2)$ Gauge Theory
The anticommutation and commutation relations for the $SU(2)$ gauge theory are
$\{\xi_{\alpha}(x),\xi_{\beta}^{\dagger}(y)\}=\delta_{xy}\delta_{\alpha\beta},\qquad[U_{\alpha\beta}(x),E_{\text{L}}^{a}(y)]=\delta_{xy}T_{\alpha\gamma}^{a}U_{\gamma\beta}(x),\qquad[U_{\alpha\beta}(x),E_{\text{R}}^{a}(y)]=-\delta_{xy}U_{\alpha\gamma}(x)T_{\gamma\beta}^{a},$ (23)
where $T^{a}=\sigma^{a}/2$. We adopt the Einstein summation convention for repeated indices. Latin indices like $a,b,c,\dots$ run over $\{1,2,3\}$ for $SU(2)$ and over $\{1,2,\dots,8\}$ for $SU(3)$, while Greek indices like $\alpha,\beta,\gamma,\dots$ run over $\{1,2\}$ for $SU(2)$ and over $\{1,2,3\}$ for $SU(3)$. In addition we have $U_{11}^{\dagger}(x)=U_{22}(x)$ and $U_{21}^{\dagger}(x)=-U_{12}(x)$.
A basis for the local Hilbert space of the gauge field is given by
$|j,m_{\text{L}},m_{\text{R}}\rangle,\qquad j=0,\frac{1}{2},1,\dots,\qquad m_{\text{L}},m_{\text{R}}=-j,-j+1,\dots,j,$ (24)
in which the operators $E^{2}$, $E^{3}_{\text{L}}$ and $E^{3}_{\text{R}}$ are diagonal and give
$E^{2}|j,m_{\text{L}},m_{\text{R}}\rangle=j(j+1)|j,m_{\text{L}},m_{\text{R}}\rangle,$ (25)
$E_{\text{L}}^{3}|j,m_{\text{L}},m_{\text{R}}\rangle=m_{\text{L}}|j,m_{\text{L}},m_{\text{R}}\rangle,\qquad E_{\text{R}}^{3}|j,m_{\text{L}},m_{\text{R}}\rangle=m_{\text{R}}|j,m_{\text{L}},m_{\text{R}}\rangle.$ (26)
The operators $U_{\alpha\beta}$ can be expressed as
$U_{\alpha\beta}|j,m_{\text{L}},m_{\text{R}}\rangle=\\ c_{\alpha\beta}^{-}(j,m_{\text{L}},m_{\text{R}})|j-\frac{1}{2},m_{\text{L}}+\frac{3}{2}-\alpha,m_{\text{R}}+\frac{3}{2}-\beta\rangle+c_{\alpha\beta}^{+}(j,m_{\text{L}},m_{\text{R}})|j+\frac{1}{2},m_{\text{L}}+\frac{3}{2}-\alpha,m_{\text{R}}+\frac{3}{2}-\beta\rangle,$ (27)
where $c_{\alpha\beta}^{\pm}(j,m_{\text{L}},m_{\text{R}})$ are products of Clebsch-Gordan coefficients,
$c_{\alpha\beta}^{\pm}(j,m_{\text{L}},m_{\text{R}})=\sqrt{\frac{2j+1}{2j\pm 1+1}}\langle j\pm\frac{1}{2},m_{\text{L}}+\frac{3}{2}-\alpha|j,m_{\text{L}};\frac{1}{2},\frac{3}{2}-\alpha\rangle\langle j\pm\frac{1}{2},m_{\text{R}}+\frac{3}{2}-\beta|j,m_{\text{R}};\frac{1}{2},\frac{3}{2}-\beta\rangle.$ (28)
We also have the decomposition
$U_{\alpha\beta}=M_{\alpha}^{\text{L}}M_{\beta}^{\text{R}}\big{(}J^{-}c_{\alpha\beta}^{-}+J^{+}c_{\alpha\beta}^{+}\big{)},\qquad\text{(no summation over {\hbox{\alpha}} and {\hbox{\beta}})}$ (29)
with
$M_{\alpha}^{\text{L}}|j,m_{\text{L}},m_{\text{R}}\rangle$ $=|j,m_{\text{L}}+\frac{3}{2}-\alpha,m_{\text{R}}\rangle,$ (30)
$M_{\beta}^{\text{R}}|j,m_{\text{L}},m_{\text{R}}\rangle$ $=|j,m_{\text{L}},m_{\text{R}}+\frac{3}{2}-\beta\rangle,$ (31)
$J^{\pm}|j,m_{\text{L}},m_{\text{R}}\rangle$ $=|j\pm\frac{1}{2},m_{\text{L}},m_{\text{R}}\rangle,$ (32)
$c_{\alpha\beta}^{\pm}|j,m_{\text{L}},m_{\text{R}}\rangle$ $=c_{\alpha\beta}^{\pm}(j,m_{\text{L}},m_{\text{R}})|j,m_{\text{L}},m_{\text{R}}\rangle.$ (33)
Analogously to the $U(1)$ theory, the $SU(2)$ theory has mesons
$\Phi_{1}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)\xi(x)-\xi^{\dagger}(x+1)\xi(x+1)\Big{]},$ (34)
$\Phi_{2}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)U^{\dagger}(x)\xi(x+1)+\xi^{\dagger}(x+1)U(x)\xi(x)\Big{]},$ (35)
$\Phi_{3}(x)$ $=\frac{1}{a}\Big{[}\xi^{\dagger}(x)U^{\dagger}(x)\xi(x+1)-\xi^{\dagger}(x+1)U(x)\xi(x)\Big{]}.$ (36)
In addition, it has a baryon,
$\Phi_{4}(x)=\frac{1}{a}\epsilon_{\alpha\beta}\xi_{\alpha}^{\dagger}(x)\xi_{\beta}^{\dagger}(x),$ (37)
where $\epsilon_{\alpha\beta}$ is the Levi-Civita symbol. However, the baryon is expected to coincide with a meson in the continuum limit.
III.3 $SU(3)$ Gauge Theory
The $SU(3)$ case is a more complicated version of the $SU(2)$ one. For the details see [35]. The basis states are labelled by eight quantum numbers
$|p,q,T_{\text{L}},T_{\text{L}}^{z},Y_{\text{L}},T_{\text{R}},T_{\text{R}}^{z},Y_{\text{R}}\rangle,$ (38)
$p,q=0,1,2,\dots,\qquad T_{{\text{L}}/{\text{R}}}=0,\frac{1}{2},\dots,\frac{1}{2}(p+q),$ (39)
$T_{{\text{L}}/{\text{R}}}^{z}=-\frac{1}{2}(p+q),-\frac{1}{2}(p+q+1),\dots,\frac{1}{2}(p+q),\qquad Y_{{\text{L}}/{\text{R}}}=-\frac{1}{3}(q+2p),-\frac{1}{3}(q+2p+1),\dots,\frac{1}{3}(p+2q).$ (40)
Similarly to the $SU(2)$ case, in this basis we have the decomposition
$U_{\alpha\beta}=M_{\alpha}^{\text{L}}M_{\beta}^{\text{R}}\big{(}P^{+}\mathcal{C}_{1\alpha}^{\text{L}}\mathcal{C}_{1\beta}^{\text{R}}\hat{N}_{1}+P^{-}Q^{+}\mathcal{C}_{2\alpha}^{\text{L}}\mathcal{C}_{2\beta}^{\text{R}}\hat{N}_{2}+Q^{-}\mathcal{C}_{3\alpha}^{\text{L}}\mathcal{C}_{3\beta}^{\text{R}}\hat{N}_{3}\big{)},\qquad\text{(no summation over {\hbox{\alpha}} and {\hbox{\beta}})}$ (41)
where ($i=\text{L},\text{R}$)
$M_{1}^{i}|T_{i}^{z},Y_{i}\rangle=|T_{i}^{z}+\frac{1}{2},Y_{i}+\frac{1}{3}\rangle,\qquad M_{2}^{i}|T_{i}^{z},Y_{i}\rangle=|T_{i}^{z}-\frac{1}{2},Y_{i}+\frac{1}{3}\rangle,\qquad M_{3}^{i}|T_{i}^{z},Y_{i}\rangle=|T_{i}^{z},Y_{i}-\frac{2}{3}\rangle,$ (42)
$P^{\pm}|p,q\rangle=|p\pm 1,q\rangle,\qquad Q^{\pm}|p,q\rangle=|p,q\pm 1\rangle,\qquad T_{i}^{\pm}|T_{i}\rangle=|T_{i}\pm\frac{1}{2}\rangle,$ (43)
$\mathcal{C}_{\alpha 1}^{i}=T_{i}^{+}\hat{C}_{\alpha 1}^{i(a)}+T_{i}^{-}\hat{C}_{\alpha 1}^{i(b)},\qquad\mathcal{C}_{\alpha 2}^{i}=T_{i}^{+}\hat{C}_{\alpha 2}^{i(a)}+T_{i}^{-}\hat{C}_{\alpha 2}^{i(b)},\qquad\mathcal{C}_{\alpha 3}^{i}=\hat{C}_{\alpha 3}^{i},$ (44)
and the diagonal operators $\hat{N}_{1,2,3},\,\hat{C}_{\alpha\beta}^{i(a,b)}$ and $\hat{C}_{\alpha 3}^{i}$ can be found in [35].
As for the $SU(2)$ case, we can suppose that the $SU(3)$ theory has mesons of the form of $\Phi_{i}$ in equations (34)-(36), but baryons are actually fermionic particles in this case,
$\Phi_{4}(x)=\frac{1}{a^{3/2}}\epsilon_{\alpha\beta\gamma}\xi_{\alpha}^{\dagger}(x)\xi_{\beta}^{\dagger}(x)\xi_{\gamma}^{\dagger}(x).$ (45)
IV Interpolating Operators on Qubits
The theories and the operators discussed in the previous section need to be mapped to a qubit system. In one dimension, the Jordan-Wigner transformation works well to map fermionic operators to qubit operators. More specifically we choose the following orderings:
$\xi(x)\rightarrow\sigma^{z}(0)\cdots\sigma^{z}(x-1)\sigma^{+}(x)$ for $U(1)$, (46)
[…]
where $\sigma^{\pm}=(\sigma^{x}\pm i\sigma^{y})/2$.
On the other hand, the local Hilbert space of the gauge field needs to be truncated. For $U(1)$ this is quite simple: if we use $k$ qubits for a single link, then we impose the cut-off $\Lambda=2^{k-1}$ and we encode the states $|-\Lambda\rangle,|-\Lambda+1\rangle,\dots,|\Lambda-1\rangle$ in the computational basis. The operator $U$ needs to be adapted to the truncation. A common choice is to take
$U|\mathcal{E}\rangle=|\mathcal{E}-1\rangle,\qquad\text{for }\mathcal{E}=-\Lambda+1,\dots,\Lambda-1,$ (49)
$U|-\Lambda\rangle=|\Lambda-1\rangle.$
In this way $U$ is a permutation operator and can be implemented on a qubit system using $O(k^{3})$ gates and no ancillary qubit, or using $O(k^{2})$ gates and $k-1$ ancillary qubits. For $SU(2)$ and $SU(3)$ the situation is more complicated but similar in essence. We refer the reader to [35] for the details. After imposing a cut-off $\Lambda$, the permutation operators $M_{\alpha}^{\text{L}/\text{R}},\,J^{\pm}$ of $SU(2)$, and $M_{\alpha}^{\text{L}/\text{R}},\,P^{\pm},\,Q^{\pm},\,T_{\text{L}/\text{R}}^{\pm}$ of $SU(3)$, take the same form as $U$ in equation (49).
The diagonal operators $c_{\alpha\beta}^{\pm}$ in (29) and the ones in (41), like for instance $\hat{C}_{11}^{\text{L}(a)}\hat{C}_{11}^{\text{R}(a)}\hat{N}_{1}$, are not unitary. However, their entries lie between $-1$ and $1$, and are efficiently computable classically. We can implement them in the following way: suppose that the function $f(x)$, where $x$ can take $M$ values encoded in the computational basis, is efficiently computable with classical techniques, and take an ancillary register of $\eta=O(\log_{2}M)$ qubits initialized in the state
$|F\rangle=\frac{1}{\sqrt{N}}\sum_{n=0}^{N-1}e^{-i2\pi n/N}|n\rangle,\qquad N=2^{\eta}.$ (50)
Then we can build efficient quantum circuits to implement unitary operators $U_{f}^{\pm}$ such that
$U_{f}^{\pm}|x\rangle|F\rangle=e^{\pm if(x)}|x\rangle|F\rangle.$ (51)
We can use LCU to combine the two operators and obtain
$\frac{(U_{f}^{+}-U_{f}^{-})}{2i}|x\rangle|F\rangle=\sin[f(x)]|x\rangle|F\rangle.$ (52)
Going back to our case, denote by $\mathcal{D}$ a diagonal operator, and by $\mathcal{D}(x)$ its entries, with $x=(j,m_{\text{L}},m_{\text{R}})$ for $SU(2)$, and $x=(p,q,T_{\text{L}},T_{\text{L}}^{z},Y_{\text{L}},T_{\text{R}},T_{\text{R}}^{z},Y_{\text{R}})$ for $SU(3)$. Then, taking $f(x)=\arcsin[\mathcal{D}(x)]$, we can implement $\mathcal{D}$. In this way we can write $U_{\alpha\beta}$ as a linear combination of unitary operators, and we can use LCU to implement it.
We can now write the interpolating operators in the qubit system. For $U(1)$ they take the form
$a\Phi_{1}(x)$ $=\sigma^{+}(x)\sigma^{-}(x)-\sigma^{+}(x+1)\sigma^{-}(x+1),$ (53)
$a\Phi_{2}(x)$ $=\sigma^{-}(x)U^{\dagger}(x)\sigma^{+}(x+1)+\sigma^{-}(x+1)U(x)\sigma^{+}(x),$ (54)
$a\Phi_{3}(x)$ $=\sigma^{-}(x)U^{\dagger}(x)\sigma^{+}(x+1)-\sigma^{-}(x+1)U(x)\sigma^{+}(x).$ (55)
For $SU(2)$ they take the form
$a\Phi_{1}(x)$ $=\sum_{\alpha}\sigma_{\alpha}^{+}(x)\sigma_{\alpha}^{-}(x)-\sigma_{\alpha}^{+}(x+1)\sigma_{\alpha}^{-}(x+1),$ (56)
$a\Phi_{2}(x)$ $=\sum_{\alpha,\beta}\sigma_{\alpha}^{-}(x)\zeta_{\alpha\beta}(x)U_{\alpha\beta}^{\dagger}(x)\sigma_{\beta}^{+}(x+1)+\sigma_{\alpha}^{-}(x+1)\zeta_{\alpha\beta}(x)U_{\alpha\beta}(x)\sigma_{\beta}^{+}(x),$ (57)
$a\Phi_{3}(x)$ $=\sum_{\alpha,\beta}\sigma_{\alpha}^{-}(x)\zeta_{\alpha\beta}(x)U_{\alpha\beta}^{\dagger}(x)\sigma_{\beta}^{+}(x+1)-\sigma_{\alpha}^{-}(x+1)\zeta_{\alpha\beta}(x)U_{\alpha\beta}(x)\sigma_{\beta}^{+}(x),$ (58)
$a\Phi_{4}(x)$ $=2\sigma_{1}^{-}(x)\sigma_{2}^{-}(x),$ (59)
where
$\zeta_{11}(x)$ $=\sigma_{2}^{z}(x),$ $\zeta_{12}(x)$ $=\sigma_{2}^{z}(x)\sigma_{1}^{z}(x+1),$ (60)
$\zeta_{21}(x)$ $=\mathbb{1},$ $\zeta_{22}(x)$ $=\sigma_{1}^{z}(x+1).$ (61)
For $SU(3)$ we have
$a\Phi_{1}(x)$ $=\sum_{\alpha}\sigma_{\alpha}^{+}(x)\sigma_{\alpha}^{-}(x)-\sigma_{\alpha}^{+}(x+1)\sigma_{\alpha}^{-}(x+1),$ (62)
$a\Phi_{2}(x)$ $=\sum_{\alpha,\beta}\sigma_{\alpha}^{-}(x)\zeta_{\alpha\beta}(x)U_{\alpha\beta}^{\dagger}(x)\sigma_{\beta}^{+}(x+1)+\sigma_{\alpha}^{-}(x+1)\zeta_{\alpha\beta}(x)U_{\alpha\beta}(x)\sigma_{\beta}^{+}(x),$ (63)
$a\Phi_{3}(x)$ $=\sum_{\alpha,\beta}\sigma_{\alpha}^{-}(x)\zeta_{\alpha\beta}(x)U_{\alpha\beta}^{\dagger}(x)\sigma_{\beta}^{+}(x+1)-\sigma_{\alpha}^{-}(x+1)\zeta_{\alpha\beta}(x)U_{\alpha\beta}(x)\sigma_{\beta}^{+}(x),$ (64)
$a^{3/2}\Phi_{4}(x)$ $=6\sigma_{1}^{z}(0)\sigma_{2}^{z}(0)\sigma_{3}^{z}(0)\sigma_{1}^{z}(1)\cdots\sigma_{3}^{z}(x-1)\sigma_{1}^{-}(x)\sigma_{2}^{-}(x)\sigma_{3}^{-}(x),$ (65)
where again $\zeta_{\alpha\beta}(x)$ are strings of $Z$ Pauli matrices analogous to the ones of the $SU(2)$ case. Notice that the baryon is a fermionic particle, which implies that the string of Pauli matrices $\sigma_{1}^{z}(0)\sigma_{2}^{z}(0)\sigma_{3}^{z}(0)\sigma_{1}^{z}(1)\cdots\sigma_{3}^{z}(x-1)$ does not cancel out as for bosonic particles. All these operators can be decomposed as a linear combination of unitary operators, and we can implement them with LCU using $O(1)$ ancillary qubits.
V Quantum Chromodynamics
V.1 Staggered Fermions in Three Dimensions
Figure 2: Identification of the lattice fields with the continuum fields.
The staggered-fermion Hamiltonian in three dimensions reads
$H_{0}=\frac{1}{2a}\sum_{\mathbf{r},\hat{\mathbf{n}}}\big{[}\xi^{\dagger}(\mathbf{r})\xi(\mathbf{r}+\hat{\mathbf{n}})+\xi^{\dagger}(\mathbf{r}+\hat{\mathbf{n}})\xi(\mathbf{r})\big{]}\eta(\mathbf{r},\hat{\mathbf{n}})+m_{0}\sum_{\mathbf{r}}(-1)^{x+y+z}\xi^{\dagger}(\mathbf{r})\xi(\mathbf{r})$ (66)
where $\mathbf{r}=(x,y,z)$ is a triplet of integers, $\hat{\mathbf{n}}\in\{\hat{\mathbf{x}},\hat{\mathbf{y}},\hat{\mathbf{z}}\}$ is a unit vector along a positive Cartesian axis, and
$\eta(\mathbf{r},\hat{\mathbf{x}})=(-1)^{z},\qquad\eta(\mathbf{r},\hat{\mathbf{y}})=(-1)^{x},\qquad\eta(\mathbf{r},\hat{\mathbf{z}})=(-1)^{y}.$ (67)
This Hamiltonian in the continuum limit describes two flavours of quarks, up and down for example, with the same mass $m_{u}=m_{d}=m_{0}$. Identifying the components of the two Dirac spinors $u(\mathbf{r})$ and $d(\mathbf{r})$ with the field $\xi(\mathbf{r})$ is not trivial, and several choices can be made up to $O(a)$ differences. What we propose here is based on [36, 37], where they perform a similar identification in momentum space. First modify $\xi(\mathbf{r})$ by a phase,
$\phi(\mathbf{r})=i^{-x-z}A(y)D(x,z)\xi(\mathbf{r}),$ (68)
with
$A(y)=\frac{1}{\sqrt{2}}\big{[}i^{y-1/2}+(-i)^{y-1/2}\big{]},\qquad D(x,z)=\frac{1}{2}\big{[}(-1)^{x}+(-1)^{z}+(-1)^{x+z+1}+1\big{]}.$ (69)
Then we identify $\phi$ at different sites with eight different fields as in figure 2,
$\begin{gathered}f_{1}(\mathbf{r})=\phi(2\mathbf{r}),\quad f_{2}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{z}}),\\ f_{3}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{z}}),\quad f_{4}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}})\\ g_{1}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}}),\quad g_{2}(\mathbf{r})=\phi(2\mathbf{r}+2\hat{\mathbf{x}}+\hat{\mathbf{y}}+\hat{\mathbf{z}}),\\ g_{3}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}}+\hat{\mathbf{z}}),\quad g_{4}(\mathbf{r})=\phi(2\mathbf{r}+2\hat{\mathbf{x}}+\hat{\mathbf{y}}).\end{gathered}$ (70)
The two Dirac spinors are obtained with the transformations
$\tilde{u}=\frac{1}{\sqrt{2}}(f+g),\qquad\tilde{d}=\frac{M}{\sqrt{2}}(f-g),$ (71)
$u=\frac{1}{\sqrt{2}}(\tilde{u}-i\tilde{d}),\qquad d=\frac{1}{\sqrt{2}}(-i\tilde{u}+\tilde{d}),$ (72)
where
$M=i\begin{pmatrix}\sigma_{y}&0\\ 0&-\sigma_{y}\end{pmatrix}=\begin{pmatrix}0&1&0&0\\ -1&0&0&0\\ 0&0&0&-1\\ 0&0&1&0\end{pmatrix}.$ (73)
We can put together the transformations as in
$u=\frac{1}{2}\big{[}(\mathbb{1}-iM)f+(\mathbb{1}+iM)g\big{]},\qquad d=\frac{1}{2}\big{[}(M-i\mathbb{1})f-(M+i\mathbb{1})g\big{]}.$ (74)
In terms of the fields $u(\mathbf{r})$ and $d(\mathbf{r})$, the free fermion Hamiltonian reads
$\begin{gathered}H_{0}=\sum_{\mathbf{r}}\bigg{[}\frac{i}{2a}\sum_{\hat{\mathbf{n}}}\Big{[}u^{\dagger}(\mathbf{r})\alpha_{n}\Delta_{\hat{\mathbf{n}}}u(\mathbf{r})+d^{\dagger}(\mathbf{r})\alpha_{n}\Delta_{\hat{\mathbf{n}}}d(\mathbf{r})\Big{]}+\frac{i}{2a}\Big{[}u^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{x}}\hat{\mathbf{y}}}u(\mathbf{r})-d^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{x}}\hat{\mathbf{y}}}d(\mathbf{r})\Big{]}\\ +m_{0}\Big{[}u^{\dagger}(\mathbf{r})\beta u(\mathbf{r})+d^{\dagger}(\mathbf{r})\beta d(\mathbf{r})\Big{]}\bigg{]}\end{gathered}$ (75)
where $\beta,\,\alpha_{n}$ and $\gamma_{5}$ are the usual Dirac matrices in the Dirac representation,
$\beta=\begin{pmatrix}\mathbb{1}&0\\ 0&-\mathbb{1}\end{pmatrix},\quad\alpha_{n}=\begin{pmatrix}0&\sigma_{n}\\ \sigma_{n}&0\end{pmatrix},\quad\gamma_{5}=\begin{pmatrix}0&\mathbb{1}\\ \mathbb{1}&0\end{pmatrix},$ (76)
and
$\begin{gathered}\Delta_{\hat{\mathbf{x}}}q(\mathbf{r})=\begin{pmatrix}q_{1}(\mathbf{r}+\hat{\mathbf{x}})-q_{1}(\mathbf{r})\\ q_{2}(\mathbf{r})-q_{2}(\mathbf{r}-\hat{\mathbf{x}})\\ q_{3}(\mathbf{r}+\hat{\mathbf{x}})-q_{3}(\mathbf{r})\\ q_{4}(\mathbf{r})-q_{4}(\mathbf{r}-\hat{\mathbf{x}})\end{pmatrix},\quad\Delta_{\hat{\mathbf{y}}}q(\mathbf{r})=\frac{1}{2}\begin{pmatrix}q_{1}(\mathbf{r})-q_{1}(\mathbf{r}-\hat{\mathbf{y}})+q_{1}(\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}})-q_{1}(\mathbf{r}+\hat{\mathbf{x}})\\ q_{2}(\mathbf{r}+\hat{\mathbf{y}})-q_{2}(\mathbf{r})+q_{2}(\mathbf{r}-\hat{\mathbf{x}})-q_{2}(\mathbf{r}-\hat{\mathbf{x}}-\hat{\mathbf{y}})\\ q_{3}(\mathbf{r})-q_{3}(\mathbf{r}-\hat{\mathbf{y}})+q_{3}(\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}})-q_{3}(\mathbf{r}+\hat{\mathbf{x}})\\ q_{4}(\mathbf{r}-\hat{\mathbf{x}})-q_{4}(\mathbf{r}-\hat{\mathbf{x}}-\hat{\mathbf{y}})+q_{4}(\mathbf{r}+\hat{\mathbf{y}})-q_{4}(\mathbf{r})\end{pmatrix},\\ \Delta_{\hat{\mathbf{z}}}q(\mathbf{r})=\begin{pmatrix}q_{1}(\mathbf{r}+\hat{\mathbf{z}})-q_{1}(\mathbf{r})\\ q_{2}(\mathbf{r})-q_{2}(\mathbf{r}-\hat{\mathbf{z}})\\ q_{3}(\mathbf{r})-q_{3}(\mathbf{r}-\hat{\mathbf{z}})\\ q_{4}(\mathbf{r}+\hat{\mathbf{z}})-q_{4}(\mathbf{r})\end{pmatrix},\quad\Delta_{\hat{\mathbf{x}}\hat{\mathbf{y}}}q(\mathbf{r})=\frac{1}{2}\begin{pmatrix}q_{1}(\mathbf{r})-q_{1}(\mathbf{r}-\hat{\mathbf{y}})-q_{1}(\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}})+q_{1}(\mathbf{r}+\hat{\mathbf{x}})\\ q_{2}(\mathbf{r}+\hat{\mathbf{y}})-q_{2}(\mathbf{r})-q_{2}(\mathbf{r}-\hat{\mathbf{x}})+q_{2}(\mathbf{r}-\hat{\mathbf{x}}-\hat{\mathbf{y}})\\ q_{3}(\mathbf{r})-q_{3}(\mathbf{r}-\hat{\mathbf{y}})-q_{3}(\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}})+q_{3}(\mathbf{r}+\hat{\mathbf{x}})\\ q_{4}(\mathbf{r}-\hat{\mathbf{x}})-q_{4}(\mathbf{r}-\hat{\mathbf{x}}-\hat{\mathbf{y}})-q_{4}(\mathbf{r}+\hat{\mathbf{y}})+q_{4}(\mathbf{r})\end{pmatrix}.\end{gathered}$ (77)
Assuming we can expand the fields in Taylor series, in the continuum limit we have
$\begin{gathered}\Delta_{\hat{\mathbf{x}}}q(\mathbf{r})=a\partial_{x}q(\mathbf{r})+O(a^{2}),\quad\Delta_{\hat{\mathbf{y}}}q(\mathbf{r})=a\partial_{y}q(\mathbf{r})+O(a^{2}),\quad\Delta_{\hat{\mathbf{z}}}q(\mathbf{r})=a\partial_{z}q(\mathbf{r})+O(a^{2}),\\ \Delta_{\hat{\mathbf{x}}\hat{\mathbf{y}}}q(\mathbf{r})=O(a^{2}),\end{gathered}$ (78)
which confirms that the free staggered-fermion Hamiltonian describes two species of Dirac fermions in the continuum limit.
Figure 3: Another identification of the lattice fields with the continuum fields.
Another possible identification of the spinor components is the following. Based on figure 3, we set
$\begin{gathered}f_{1}(\mathbf{r})=\phi(2\mathbf{r}),\quad f_{2}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{z}}),\quad f_{3}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{z}}),\quad f_{4}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}})\\ g_{1}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}}),\quad g_{2}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{y}}+\hat{\mathbf{z}}),\quad g_{3}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}}+\hat{\mathbf{z}}),\quad g_{4}(\mathbf{r})=\phi(2\mathbf{r}+\hat{\mathbf{y}}),\end{gathered}$ (79)
and
$u=\frac{1}{\sqrt{2}}(f+g),\qquad d=\frac{M}{\sqrt{2}}(f-g).$ (80)
Then the Hamiltonian takes the form
$\begin{gathered}H_{0}=\sum_{\mathbf{r}}\bigg{[}\frac{i}{2a}\sum_{\hat{\mathbf{n}}}\Big{[}u^{\dagger}(\mathbf{r})\alpha_{n}\Delta_{\hat{\mathbf{n}}}u(\mathbf{r})+d^{\dagger}(\mathbf{r})\alpha_{n}\Delta_{\hat{\mathbf{n}}}d(\mathbf{r})\Big{]}\\ +\frac{i}{2a}\Big{[}u^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{x}}}^{2}d(\mathbf{r})+d^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{x}}}^{2}u(\mathbf{r})+iu^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{y}}}^{2}d(\mathbf{r})-id^{\dagger}(\mathbf{r})\beta\gamma_{5}\Delta_{\hat{\mathbf{y}}}^{2}u(\mathbf{r})\Big{]}\\ +m_{0}\Big{[}u^{\dagger}(\mathbf{r})\beta u(\mathbf{r})+d^{\dagger}(\mathbf{r})\beta d(\mathbf{r})\Big{]}\bigg{]},\end{gathered}$ (81)
where now
$\Delta_{\hat{\mathbf{x}}}q(\mathbf{r})$ $=\frac{1}{2}\big{[}q(\mathbf{r}+\hat{\mathbf{x}})-q(\mathbf{r}-\hat{\mathbf{x}})\big{]},$ $\qquad\Delta_{\hat{\mathbf{x}}}^{2}q(\mathbf{r})$ $=\frac{1}{2}\big{[}q(\mathbf{r}+\hat{\mathbf{x}})+q(\mathbf{r}-\hat{\mathbf{x}})-2q(\mathbf{r})\big{]},$ (82)
$\Delta_{\hat{\mathbf{y}}}q(\mathbf{r})$ $=\frac{1}{2}\big{[}q(\mathbf{r}+\hat{\mathbf{y}})-q(\mathbf{r}-\hat{\mathbf{y}})\big{]},$ $\qquad\Delta_{\hat{\mathbf{y}}}^{2}q(\mathbf{r})$ $=\frac{1}{2}\big{[}q(\mathbf{r}+\hat{\mathbf{y}})+q(\mathbf{r}-\hat{\mathbf{y}})-2q(\mathbf{r})\big{]},$
$\Delta_{\hat{\mathbf{z}}}q(\mathbf{r})$ $=\begin{pmatrix}q_{1}(\mathbf{r}+\hat{\mathbf{z}})-q_{1}(\mathbf{r})\\ q_{2}(\mathbf{r})-q_{2}(\mathbf{r}-\hat{\mathbf{z}})\\ q_{3}(\mathbf{r})-q_{3}(\mathbf{r}-\hat{\mathbf{z}})\\ q_{4}(\mathbf{r}+\hat{\mathbf{z}})-q_{4}(\mathbf{r})\end{pmatrix}.$
The main difference between the two identifications is that in the second one there are mixing terms between the two flavors already in the free theory. For this reason, in the following we consider only the first identification, equations (70) and (74).
V.2 Interpolators in the Kogut-Susskind Formulation
The Kogut-Susskind Hamiltonian for quantum chromodynamics on a three-dimensional cubic lattice is
$H$ $=H_{\text{M}}+H_{\text{K}}+H_{\text{E}}+H_{\text{B}},$ (83)
$H_{\text{M}}$ $=m_{0}\sum_{\mathbf{r}}(-1)^{x+y+z}\xi^{\dagger}(\mathbf{r})\xi(\mathbf{r}),$ (84)
$H_{\text{K}}$ $=\frac{1}{2a}\sum_{\mathbf{r},\hat{\mathbf{n}}}\big{[}\xi^{\dagger}(\mathbf{r})U(\mathbf{r},\hat{\mathbf{n}})\xi(\mathbf{r}+\hat{\mathbf{n}})+\xi^{\dagger}(\mathbf{r}+\hat{\mathbf{n}})U^{\dagger}(\mathbf{r},\hat{\mathbf{n}})\xi(\mathbf{r})\big{]}\eta(\mathbf{r},\hat{\mathbf{n}}),$ (85)
$H_{\text{E}}$ $=\frac{g_{0}^{2}}{2a}\sum_{\mathbf{r},\hat{\mathbf{n}}}E(\mathbf{r},\hat{\mathbf{n}})^{2},$ (86)
[…]
where $P_{\square}$ is a plaquette operator.
Our goal here is to find interpolating operators for the particles appearing in QCD, like the pion and the nucleons. Knowing the correspondence between the fields $u$ and $d$ and $\xi$, equations (70) and (74), we can use results from the literature on traditional lattice calculations [38, 39, 40, 41]. However, the symmetries of the formulation used in [38, 39, 40, 41] are quite different from the symmetries of the Kogut-Susskind Hamiltonian. In particular, cubic and flavor symmetries are intertwined between each other for staggered fermions, which implies that spin and isospin mix together. On the other hand, spin and isospin are among the main symmetries used to classify and identify particles. As a consequence, identification of particles in the Kogut-Susskind formulation is possible only near the continuum limit, and up to $O(a)$ errors.
As examples, consider the pions and the nucleons. Pion interpolators are
$\begin{gathered}\hat{\mathcal{O}}_{\pi^{+}}(\mathbf{r})=d^{\dagger}(\mathbf{r})\gamma_{0}\gamma_{5}u(\mathbf{r}),\qquad\hat{\mathcal{O}}_{\pi^{-}}(\mathbf{r})=u^{\dagger}(\mathbf{r})\gamma_{0}\gamma_{5}d(\mathbf{r}),\\ \hat{\mathcal{O}}_{\pi^{0}}(\mathbf{r})=\frac{1}{\sqrt{2}}\big{[}u^{\dagger}(\mathbf{r})\gamma_{0}\gamma_{5}u(\mathbf{r})-d^{\dagger}(\mathbf{r})\gamma_{0}\gamma_{5}d(\mathbf{r})\big{]}.\end{gathered}$ (88)
A proton interpolator is
$\hat{\mathcal{O}}_{p}(\mathbf{r})=\epsilon_{abc}\big{[}u_{a}^{T}(\mathbf{r})C\gamma_{5}d_{b}(\mathbf{r})\big{]}u_{c}(\mathbf{r}),$ (89)
where $C=-i\alpha_{y}$ is the charge-conjugation matrix in the Dirac representation. The neutron interpolator is obtained by exchanging $u$ and $d$.
These operators are not gauge invariant in the Kogut-Susskind formulation, because they involve products between $\xi$ operators at different lattice sites as a consequence of the identifications in equation (70). We need to insert appropriate link operators to make them gauge invariant. The mesonic operators contain terms like $\xi^{\dagger}(2\mathbf{r})\xi(2\mathbf{r}+\hat{\mathbf{x}})$, which is made gauge invariant by $\xi^{\dagger}(2\mathbf{r})U(2\mathbf{r},\hat{\mathbf{x}})\xi(2\mathbf{r}+\hat{\mathbf{x}})$. The baryonic operators contain terms like $\epsilon_{abc}\xi_{a}(2\mathbf{r})\xi_{b}(2\mathbf{r}+\hat{\mathbf{x}})\xi_{c}(2\mathbf{r}+2\hat{\mathbf{x}}+\hat{\mathbf{y}})$, which is made gauge invariant by
$\epsilon_{abc}U_{aa^{\prime}}(2\mathbf{r}+\hat{\mathbf{x}},-\hat{\mathbf{x}})\xi_{a^{\prime}}(2\mathbf{r})\xi_{b}(2\mathbf{r}+\hat{\mathbf{x}})U_{cc^{\prime}}(2\mathbf{r}+\hat{\mathbf{x}},\hat{\mathbf{y}})U_{c^{\prime}c^{\prime\prime}}(2\mathbf{r}+\hat{\mathbf{x}}+\hat{\mathbf{y}},\hat{\mathbf{x}})\xi_{c^{\prime}}(2\mathbf{r}+2\hat{\mathbf{x}}+\hat{\mathbf{y}}).$ (90)
Following the examples given in the previous section on one-dimensional lattices, we can map the interpolating operators into qubit operators, and decompose them using LCU. The link operators $U_{\alpha\beta}$ take the same form as in the one dimensional case. The main difference with the previous section is in the encoding of the fermionic field $\xi$. According to [42], it is computationally more convenient to use the generalized superfast encoding (GSE) instead of a Jordan-Wigner transformation in three dimensions. We review the GSE in the appendix, where we also explain how to encode odd fermionic operators that are necessary to implement the nucleon interpolators.
We conclude this section with some considerations on the success probability $\rho$ in equation (7). The factor $C$ depends on the number of terms in the interpolating operator and on its mass dimension. Each quark field contains a factor $1/a^{1.5}$ in the continuum normalization. Counting also the numerical coefficients in the expansions of the interpolators in terms of $\xi$, we obtain approximately $C_{\pi^{\pm}}\sim 432/a^{3}$, $C_{\pi^{0}}=\sqrt{2}C_{\pi^{\pm}}$, and $C_{p}\sim 11500/a^{4.5}$. The large numerical coefficients appearing here are a consequence of choosing Kogut-Susskind fermions. For a comparison, the corresponding coefficients using Wilson fermions are $C_{\pi^{\pm}}=12/a^{3}$, $C_{\pi^{0}}=\sqrt{2}C_{\pi^{\pm}}$, and $C_{p}=24/a^{4.5}$.
VI Conclusions
In reference [25] we showed how to create wavepackets from the vacuum in scalar field theories. Here we focused on lattice gauge theories, and increased gradually the level of difficulty, starting from a $U(1)$ theory in one dimension and ending with QCD in three dimensions with two flavors of quarks. We considered different kind of particles, namely pions and nucleons in QCD.
Three ingredients are necessary to apply our method. First, we need to prepare the vacuum of the theory. This can be done with variational methods, like [43, 24], or with an adiabatic transformation [44]. Second, we need to know the spectrum of the theory. Near the continuum limit, this amounts to know the particle content and the particle masses. Finally, we need interpolating operators for the particles we aim at. These operators are built on a symmetry basis, and they should carry the quantum numbers of the particles. Our work is further extensible to other theories, formulations and particles, with the only requirement that there has to be a mass gap.
The creation of a wavepacket with our strategy requires the use of LCU. This technique succeeds with a probability $\rho$ that depends on the theory, on the wavepacket, and on the particle. If we want to prepare $n$ wavepackets, the initial state preparations needs to be repeated on average $O(1/\rho^{n})$ times to obtain the correct state for the simulation, so we can create only a small number of incoming wavepackets. Fortunately, $n=2$ is the most relevant scenario. The success probability seems to be the main limitation of our approach, but at least it does not affect the circuit depth, which is a favorable feature.
We chose to work in the Kogut-Susskind framework. For QCD, we imported interpolators for the pions and the nucleons from the literature on traditional lattice calculations. We transformed these operators into operators in the Kogut-Susskind formulation, but this rises some problems. There are many possible choices for these transformations, with differences of the order of the lattice spacing, and none of them respects the symmetries that determine the particles in the continuum. As a consequence, particle identification is more problematic, a well-known downside of staggered fermions. Furthermore, the success probability suffers significantly from the transformation of staggered fermions into Dirac spinors. The interpolating operators in the Kogut-Susskind formulation become long linear expansions of operators, with insertion of gauge fields to achieve gauge invariance. The resulting overheads in the success probability are $\sim 10^{4}$ for the pions, and $\sim 10^{8}$ for the nucleons. If we use Wilson fermions, the equivalent overheads are $\sim 10^{2}$.
In the future, we aim to provide more precise estimates on the resources needed, especially in the cases that have already been tested [23, 24] to make a comparison between the different strategies. If the hardware is mature enough, it would be useful to test the method on a real quantum computer. It would also be important to have asymptotical estimates of the other approaches in the literature to determine which one is the most promising.
VII Acknowledgments
We thank the support from FCT – Fundação para a Ciência e a Tecnologia (Portugal), namely through project UIDB/04540/2020, as well as from projects QuantHEP and HQCC supported by the EU QuantERA ERA-NET Cofund in Quantum Technologies and by FCT (QuantERA/004/2021). Finally, Matteo Turco thanks FCT for support through the grant PRT/BD/154668/2022.
Appendix A Generalised Superfast Encoding of Fermions
We review the generalised superfast encoding (GSE) of fermions into qubits [45]. In [42], this encoding was used to estimate the resource requirements for quantum simulation of lattice gauge theories such as quantum chromodynamics, showing a substantial improvement with respect to a Jordan-Wigner transformation. As presented in [45], the GSE allows for the representation of even fermionic operators only. As odd operators are necessary for the preparation of fermionic-particle wavepackets, here we also show how they can be represented in the GSE. In this appendix we drop the Einstein convention for repeated indices.
A.1 Fermions
We first introduce some general facts about fermions. For a more detailed discussion see also [46]. Consider a system of $N$ independent fermionic modes with annihilation operators $a_{1},\dots,a_{N}$ satisfying the usual anticommutation relations
$\{a_{I},a_{J}\}=0,\qquad\{a_{I},a_{J}^{\dagger}\}=\delta_{IJ}.$ (91)
It is useful to introduce the Majorana modes as
$\gamma_{I}^{+}=a_{I}+a_{I}^{\dagger},\qquad\gamma_{I}^{-}=-i(a_{I}-a_{I}^{\dagger}).$ (92)
The Majorana modes are Hermitian, ${\gamma_{I}^{\pm}}^{\dagger}=\gamma_{I}^{\pm}$, and satisfy the anticommutation relations
$\{\gamma_{I}^{\rho},\gamma_{J}^{\sigma}\}=2\delta_{IJ}\delta_{\rho\sigma},$ (93)
where $\rho,\sigma\in\{+,-\}$. In particular, note that ${\gamma_{I}^{\pm}}^{2}=1$.
The Hilbert space of a fermionic system is spanned by all possible polynomials in the fermionic operators applied to the vacuum, and has dimension $2^{N}$. The vacuum is defined as the state annihilated by all the annihilation operators,
$a_{I}|0\rangle=0,\quad\forall I.$ (94)
There are two subspaces of dimension $2^{N-1}$ each, usually called the even and the odd subspace. The even subspace is spanned by arbitrary combinations of products with an even number of fermionic operators applied to the vacuum. Similarly, the odd subspace is spanned by odd products of operators. Combinations of states in different subspaces are physically forbidden.
The even operators form a subalgebra generated by
$B_{I}=-i\gamma_{I}^{+}\gamma_{I}^{-},\qquad A_{IJ}=-i\gamma_{I}^{+}\gamma_{J}^{+}.$ (95)
These operators satisfy the following relations:
$B_{I}^{\dagger}=B_{I},\qquad A_{IJ}^{\dagger}=A_{IJ},\qquad B_{I}^{2}=1,\qquad A_{IJ}^{2}=1,\qquad A_{IJ}=-A_{JI},$ (96)
$B_{I}B_{J}=B_{J}B_{I},\qquad A_{IJ}B_{K}=(-1)^{\delta_{IK}+\delta_{JK}}B_{K}A_{IJ},\qquad A_{IJ}A_{KL}=(-1)^{\delta_{IK}+\delta_{IL}+\delta_{JK}+\delta_{JL}}A_{KL}A_{IJ},$ (97)
$A(\zeta)=i^{s}A_{\zeta(1)\zeta(2)}A_{\zeta(2)\zeta(3)}\dots A_{\zeta(s)\zeta(s+1)}=1,$ (98)
where $\zeta:\,\{1,\dots,s+1\}\rightarrow\{1,\dots,N\}$ is a closed loop of length $s$ with $\zeta(s+1)=\zeta(1)$. Furthermore we have
$\prod_{I=1}^{N}B_{I}$ $=1$ on the even subspace, (99)
$\prod_{I=1}^{N}B_{I}$ $=-1$ on the odd subspace. (100)
Given an arbitrary $\gamma_{S}^{+}$, we can write any other Majorana mode as a product of $\gamma_{S}^{+}$ and an even operator:
$\gamma_{S}^{-}=i\gamma_{S}^{+}B_{S},\qquad\gamma_{I}^{+}=i\gamma_{S}^{+}A_{SI},\qquad\gamma_{I}^{-}=-\gamma_{S}^{+}A_{SI}B_{I}.$ (101)
This feature makes it very simple to implement an arbitrary fermionic operator on a qubit system once we have an encoding for the even operators.
A.2 Even Operators in the GSE
To describe the GSE, we put the $N$ fermionic modes on a connected graph with $N$ vertices. We also require that each vertex be of even degree, and we assign an orientation to each edge. The graph can be chosen freely to suite a specific problem we want to address. For example, we might consider a Hamiltonian $H$ and a graph $G=(V,E)$ such that $H$ consists of terms acting nontrivially only on pairs of nearest neighbours on $G$, namely
$H=\sum_{(I,J)\in E}H_{IJ}.$ (102)
We denote by $d(I)$ the degree of vertex $I\in V$, and we put $d(I)/2$ qubits on $I$. Next, we introduce $d(I)$ local Majorana modes for each vertex, $\gamma_{I1},\dots,\gamma_{Id(I)}$, which satisfy the relations
$\{\gamma_{Ii},\gamma_{Ij}\}$ $=2\delta{ij},$ (103)
$[\gamma_{Ii},\gamma_{Jj}]$ $=0,\qquad\text{if $I\neq J$}.$ (104)
We order the edges connected to vertex $I$ as $1,\dots,d(I)$, and we attach the mode $\gamma_{Ii}$ to the edge $i$. The local Majorana modes can be constructed with appropriate strings of Pauli matrices acting on the qubits of vertex $I$. From these modes we define the operators
$\tilde{B}_{I}$ $=(-i)^{d(I)/2}\gamma_{I1}\cdots\gamma_{Id(I)}$ for each vertex $I$, (105)
$\tilde{A}_{IJ}$ $=\epsilon_{IJ}\gamma_{Ii}\gamma_{Jj}$ for each edge $(I,J)$, (106)
where $\epsilon_{IJ}=1$ if we follow the arrow when going from $I$ to $J$, and $\epsilon_{IJ}=-1$ if we go in the opposite direction. The modes $\gamma_{Ii}$ and $\gamma_{Jj}$ in (106) are the two modes attached to the ends of edge $(I,J)$.
The operators defined in the expressions (105) and (106) satisfy the properties (96) and (97), but not the property (98), so we need to impose it. We define the qubit counterparts of the loop operators in (98) as
$\tilde{A}(\zeta)=i^{s}\tilde{A}_{\zeta(1)\zeta(2)}\tilde{A}_{\zeta(2)\zeta(3)}\dots\tilde{A}_{\zeta(s)\zeta(s+1)}$ (107)
for any closed loop $\zeta$ of length $s$ on the graph $G$, where now $\zeta(i)$ and $\zeta(i+1)$ are connected by an edge. The group $\mathcal{S}$ generated by these loop operators form an Abelian group that commutes with $\tilde{A}_{IJ}$ and $\tilde{B}_{I}$, and can be used as a stabilizer of the subspace
=$\mathcal{L}=\{|\psi\rangle:\tilde{A}(\zeta)|\psi\rangle=|\psi\rangle\text{ forall loops }\zeta\}.$ (108)
We can identify either the even or the odd subspace of the fermionic Hilbert space with the stabilized space $\mathcal{L}$, which has dimension $2^{N-1}$. The operators $\tilde{A}_{IJ}$ and $\tilde{B}_{I}$, when restricted to this subspace, satisfy also the property (98) by construction of $\mathcal{L}$, and give us a qubit representation of the algebra of even operators. Since $G$ is a connected even-degree graph, there exists an Eulerian cycle, namely a loop $\eta$ that utilizes every edge of the graph exactly once. For this loop we have
$\prod_{I=1}^{N}\tilde{B}_{I}=\pm\tilde{A}(\eta),$ (109)
where the sign depends on our choice of the orientation of the edges. If our orientation produces a positive sign, $\mathcal{L}$ represents the even subspace. Then, if we switch the direction of one arrow, we obtain a different subspace $\mathcal{L}$, which represents the odd subspace.
A.3 Odd Operators in the GSE
In this section we show how to extend the Generalized Superfast Encoding described in the previous section to also represent odd fermionic operators. Given a graph, a specific orientation of the edges gives one of many possible stabilized subspaces. Each of them can be identified either with the even or the odd subspace of the fermionic Hilbert space. Even operators, and their qubit representations, preserve these subspaces, while odd operators do not in the fermionic space and neither should their representation on the qubit space. Our goal is to find qubit operators representing the Majorana modes in (92) and (93). This representation should be compatible with the representation $\tilde{A}_{IJ}$ and $\tilde{B}_{IJ}$ of the previous section, at least on the stabilized subspace, and should make us jump from a space where equation (109) holds with the positive sign, to a space where it holds with the negative sign, and vice versa.
Consider again a graph as in the previous section, with every vertex of even degree, and take two vertices $S$ and $T$ connected by an edge. For every vertex $I$ in the graph take a path $\zeta_{SI}$ from $S$ to $I$ that does not use the edge $(S,T)$, with $\zeta_{SI}(1)=S$ and $\zeta_{SI}(s)=I$. We can always find such paths following the Eulerian cycle $\eta$. Consider now the local Majorana mode on $S$ attached to the edge $(S,T)$, which we can call $\gamma_{S1}$ (up to relabelling of the modes). Motivated by
$\gamma_{I}^{+}$ $=i\gamma_{S}^{+}A_{SI},$ (110)
$\gamma_{I}^{-}$ $=-\gamma_{S}^{+}A_{SI}B_{I},$ (111)
as a representation of the fermionic operators we take
$\gamma_{I}^{+}$ $\rightarrow\hat{\gamma}_{I}^{+}=\gamma_{S1}\tilde{A}(\zeta_{SI}),$ (112)
$\gamma_{I}^{-}$ $\rightarrow\hat{\gamma}_{I}^{-}=i\gamma_{S1}\tilde{A}(\zeta_{SI})\tilde{B}_{I},$ (113)
where
$\tilde{A}(\zeta_{SS})$ $=1,$ (114)
$\tilde{A}(\zeta_{SI})$ $=i^{s-1}\tilde{A}_{\zeta_{SI}(1)\zeta_{SI}(2)}\tilde{A}_{\zeta_{SI}(2)\zeta_{SI}(3)}\dots\tilde{A}_{\zeta_{SI}(s-1)\zeta_{SI}(s)}.$ (115)
To see that this is a good representation of $\gamma_{I}^{\pm}$, first notice that
$[\gamma_{S1},\tilde{A}_{ST}]=0,$ (116)
$\{\gamma_{S1},\tilde{A}_{SJ}\}=0\qquad\text{if $J\neq T$}.$ (117)
The difference between the first and the second line is the reason for our choice of the paths $\zeta_{SI}$. Then, from (96), (97) and (117), it is easy to see that $\hat{\gamma}_{I}^{\pm}$ satisfy the same anticommutation relations as $\gamma_{I}^{\pm}$.
The generators of the even algebra obtained from $\hat{\gamma}_{I}^{+}$ and $\hat{\gamma}_{I}^{-}$ are
$\hat{B}_{I}=-i\hat{\gamma}_{I}^{+}\hat{\gamma}_{I}^{-},\qquad\hat{A}_{IJ}=-i\hat{\gamma}_{I}^{+}\hat{\gamma}_{J}^{+},$ (118)
and it is straightforward to see that $\hat{B}_{I}=\tilde{B}_{I}$. Suppose now there is an edge $(I,J)$. In general, we have
$\hat{A}_{IJ}\neq\tilde{A}_{IJ}.$ (119)
However, on the subspace =$\mathcal{L}=\{|\psi\rangle:\tilde{A}(\zeta)|\psi\rangle=|\psi\rangle\}$, we have
$\hat{A}_{IJ}|\psi\rangle=\hat{A}_{IJ}\tilde{A}(\zeta)|\psi\rangle.$ (120)
If we choose the loop $\zeta$ as the concatenation of $\zeta_{SI}$, $(I,J)$ and the reverse of $\zeta_{SJ}$, we obtain $\hat{A}_{IJ}\tilde{A}(\zeta)=\tilde{A}_{IJ}$, and we can conclude
$\hat{A}_{IJ}|\psi\rangle=\tilde{A}_{IJ}|\psi\rangle\quad\forall|\psi\rangle\in\mathcal{L}.$ (121)
From this we can see that the representation of the even algebra $\tilde{B}_{I}$ and $\tilde{A}_{IJ}$ is embedded in many representations of $\hat{\gamma}_{I}^{+}$ and $\hat{\gamma}_{I}^{-}$, one for each possible choice of the paths $\zeta_{SI}$. When we restrict these representations to the stabilized space $\mathcal{L}$, they all coincide and give the representation of the even algebra introduced in [45].
To conclude, we need to see what happens when we act with an odd operator on a state in the stabilized subspace $\mathcal{L}$. Since all the operators $\tilde{A}_{IJ}$ and $\tilde{B}_{I}$ commute with the stabilizer operators $\tilde{A}(\zeta)$, the relations (116) and (117) imply
$\{\hat{\gamma}_{I}^{\pm},\tilde{A}(\zeta)\}=0$ if the loop $\zeta$ involves $(S,T)$; (122)
$[\hat{\gamma}_{I}^{\pm},\tilde{A}(\zeta)]=0$ if the loop $\zeta$ does not involve $(S,T)$. (123)
Suppose now that our initial choice of the arrows in the graph allows us to identify the space stabilized by $\{\tilde{A}(\zeta)\}$ with the even subspace,
+=$\mathcal{L}^{+}=\{|\psi\rangle:\tilde{A}(\zeta)|\psi\rangle=|\psi\rangle\text{for all loops }\zeta\},\qquad\prod_{I=1}^{N}\tilde{B}_{I}=\tilde{A}(\eta)$ (124)
for an Eulerian cycle $\eta$. If we define
$\bar{A}_{ST}$ $=-\tilde{A}_{ST},$ (125)
$\bar{A}_{IJ}$ $=\tilde{A}_{IJ}\qquad\text{if $\{I,J\}\neq\{S,T\}$},$ (126)
– which is equivalent to switching the arrow of $(S,T)$ – then $\{\bar{A}(\zeta)\}$ is a stabilizer for the odd subspace,
=$\mathcal{L}^{-}=\{|\psi\rangle:\bar{A}(\zeta)|\psi\rangle=|\psi\rangle\text{ for all loops }\zeta\},\qquad\prod_{I=1}^{N}\tilde{B}_{I}=-\bar{A}(\eta).$ (127)
As a consequence of the relations (122) and (123), any odd operator makes us jump from $\mathcal{L}^{+}$ to $\mathcal{L}^{-}$, and vice versa. In virtue of this, we can conclude that the operators $\hat{\gamma}_{I}^{\pm}$, together with the two spaces $\mathcal{L}^{\pm}$ provide us with a full representation of the fermionic Hilbert space and the associated algebra.
A.4 GSE for Hypercubic Lattices
In this work we use the GSE for $SU(N)$ gauge theories on hypercubic lattices with one flavour of fermions, which means that we have $N$ independent fermionic modes on each site. In the one-dimensional case, the GSE essentially reduces to the Jordan-Wigner transformation. In $d$ dimensions we use $N+d-1$ qubits per lattice site. As an example, figure 4 depicts two neighbouring sites in an $SU(3)$ theory on a square lattice. The gray lines represent physical interaction terms in the Hamiltonian, and there is no such line between modes on the same site. The thick lines are edges of the graph we use for our encoding, and we also show the associated local Majorana modes on each mode. To obtain an interaction term between two modes, we simply choose the shortest path on the graph connecting the two modes, and the resulting encoded operator is in general a string of operators acting on the qubits involved in the path. Two out of three fermionic modes use one qubit each, the other modes use two qubits each. If mode $I$ uses one qubit, we take as local Majorana modes
$\gamma_{I1}=\sigma^{x},\qquad\gamma_{I2}=\sigma^{y}.$ (128)
If mode $I$ uses two qubits, we take
$\gamma_{I1}=\sigma^{x}\otimes\mathbb{1},\qquad\gamma_{I2}=\sigma^{y}\otimes\mathbb{1},\qquad\gamma_{I3}=\sigma^{z}\otimes\sigma^{x},\qquad\gamma_{I4}=\sigma^{z}\otimes\sigma^{y}.$ (129)
Adaptation of this example to an $SU(3)$ theory on a cubic lattice is straightforward.
Figure 4: GSE graph for $SU(3)$ on a square lattice. The gray lines represent physical interaction terms in the Hamiltonian, and there is no such line between modes on the same site. The thick lines are edges of the graph we use for our encoding, and we also show the associated local Majorana modes on each mode.